OpenStudy (cashduke):

What distance would you predict for the ball after it has fallen for 10.0 seconds? velocity increases by 10 for every second constant acceleration of 10

OpenStudy (mhchen):

Assuming the ball begins with initial velocity of 0: \[x_{f}=x_{0}+v_{0}t+\frac{1}{2}at^2\] Using this formula \[x_{f}=0+0(10s)+\frac{1}{2}(10\frac{m}{s^2})(10s)^2\] Can you solve that?